OpenStudy (anonymous):

which exponent makes the statement true? 1/5^9 = 5^? 1) 9 2) -9 3) 1/9 4) -1/9

OpenStudy (anonymous):

You first have to figure out what 5^9 is and then you just divide it by 1.

OpenStudy (anonymous):

The answer would probably be around 5.12

OpenStudy (anonymous):

http://openstudy.com/study# thats not a choice though

OpenStudy (aum):

\[\large \frac{1}{a^n} = a^{-n} \]

OpenStudy (aum):

\[\large \frac{1}{5^9} = 5^{-9} \]